HanAkoSolo (jamierox4ev3r):

Someone check my work please? (ew algebra review) What is the value of x in the equation \(\frac{2}{x}\)-3=\(\frac{26}{x}\)

HanAkoSolo (jamierox4ev3r):

\(\Large\frac{2}{x}\)-3=\(\Large\frac{26}{x}\) -2/x -2/x ----------- -3=\(\Large\frac{26}{x}\)-\(\Large\frac{2}{x}\) (\(\Large\frac{1}{24}\))-3=\(\Large\frac{24}{x}\)(\(\Large\frac{1}{24}\)) x= -\(\Large\frac{1}{8}\)

jimthompson5910 (jim_thompson5910):

you should have \[\Large \frac{1}{x} = -\frac{1}{8}\]

jimthompson5910 (jim_thompson5910):

I'm sure you see how to solve for x from here

HanAkoSolo (jamierox4ev3r):

oh i see! knew something wasn't right, thanks!